OpenStudy (anonymous):

A right triangle has a hypotenuse of length 21 and a leg of length 14. What is the length of the other leg of the triangle?

OpenStudy (johnweldon1993):

Remember pythagorean theorem \[a^2 + b^2 = c^2\] c = hypotenuse a and b are the 2 other legs